# Constrained Decoding for Secure Code Generation

### v1

@@ 4.4. Non-autoregressive Decoding
Non-autoregressive decoding generates all tokens in the output sequence together. The decoding procedure first initializes output tokens, and then uses gradients of some function to update the tokens. An example function could be a language model loss function, an energy function, or some task-specific function. Non-autoregressive decoding methods have shown promising results for machine translation (Hoang et al., 2017), reasoning and counterfactual story generation (Qin et al., 2022), and generative commonsense reasoning (Kumar et al., 2021, 2022).
@@
Recent papers argue that non-autoregressive decoding is better than autoregressive decoding for the problem of controlled text generation under constraints (Kumar et al., 2021; Qin et al., 2022; Kumar et al., 2022). The same arguments hold for code generation under constraints. During autoregressive decoding, we cannot evaluate the properties of the entire program during the generation because only a partial program is available at every step. For example, if the partially generated code has not sanitized untrusted user input yet, it does not mean that the entire generated code would not sanitize untrusted user input, so we cannot know whether the partial program is safe or not safe. On the contrary, non-autoregressive decoding generates the entire program altogether, which enables us to evaluate constraints as well as enforce constraints over the whole program.
@@
To the best of our knowledge, non-autoregressive decoding has not been evaluated on code generation before, but only text generation. In particular, the state-of-the-art scheme MuCoLa (Kumar et al., 2022) has achieved strong results of constrained text generation for common sense reasoning, beating previous methods. Therefore, we study MuCoLa and adapt it for code generation.
@@
(10) $\mathbf{y}$ $\sim P(\mathbf{y}|\mathbf{x}),$
@@
MuCoLa formulates decoding as sampling from an energy-based model (EBM) using Langevin Dynamics, following the approach in COLD decoding (Qin et al., 2022). In other words, MuCoLa performs sampling by iteratively updating the embeddings of the output sequence using gradients of the energy function. They define the energy function as the following:
(11) $\mathcal{E}(\tilde{\mathbf{e}})=-\log P(\tilde{\mathbf{e}}|\mathbf{x})-\sum_{i=1}^{C}\lambda_{i}\left(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\right).$
Here, $\lambda_{i}$ is used to balance between the output fluency and satisfying constraints. MuCoLa uses gradients to perform sampling, with details in Appendix B. The gradient update procedure will converge to sampling from the energy-based distribution (Welling and Teh, 2011).
@@
We adapt MuCoLa for constrained code generation using our correctness constraints and security constraints. We can separate our constraints into positive constraints and negative constraints. Positive constraints are key phrases that we would like to appear in generated outputs, and negative constraints are key phrases we want to block, where each phrase consists of multiple tokens. We have in total $C^{+}$ positive constraints, and $C^{-}$ negative constraints.
MuCoLa provides a differentiable positive key phrase function $f$ (details in Appendix B). We use that as a building block to define our own energy function:
(12) $\mathcal{E}^{\prime}(\tilde{\mathbf{e}})=-\log P(\tilde{\mathbf{e}}|\mathbf{x})-\sum_{i=1}^{C^{+}}\lambda_{i}\left(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\right)-\sum_{j=1}^{C^{-}}\lambda_{j}\left(f_{j}(\tilde{\mathbf{e}})-\epsilon_{j}\right)$
@@ 5. Evaluation
  • RQ2. If we use our new metrics to compare a baseline Code LLM against the state-of-the-art prefix tuning defense SVEN (He and Vechev, 2023), how does that change the conclusions about the defense? (Section 5.2)
  • RQ3. Can constrained decoding improve the security and correctness of code generated by Code LLMs? Can we use constrained decoding with prefix tuning together? (Section 5.3)
@@ 5.1. Experiment Setup
We evaluate three models in total. Two of them are open-source, decoder-only pre-trained models. They are the multi-language version CodeGen-2.7B and StarCoder2-3B, state-of-the-art (SOTA) open-source Code LLMs. Both models support Python and C/C++, the main programming languages in the Copilot dataset (Pearce et al., 2022) and our CodeGuard+. In addition, we study the SOTA prefix tuning defense. We use the trained prefix on CodeGen-2.7B to generate secure code released by the authors of SVEN (He and Vechev, 2023). We refer to the secure CodeGen-2.7B model with prefix tuning as SVEN.
@@
We use the CodeGuard+ introduced in Section 3.1 as the test suite for all evaluations. This suite comprises 23 prompts covering 9 CWEs and 2 programming languages. We use our unit tests to evaluate correctness. Consistent with related works (He and Vechev, 2023; Pearce et al., 2022; Hajipour, Hossein and Hassler, Keno and Holz, Thorsten and Schönherr, Lea and Fritz, Mario, 2023), we use CodeQL to evaluate the security of generated code. We present our evaluation results using four primary metrics: SVEN-SR, pass@$1$, secure@$1_{\text{pass}}$, and secure-pass@$1$, which are defined in Section 3.2. For constrained decoding schemes, we also calculate Constraint Rate by counting the number of outputs that satisfy constraints over the number of total outputs generated.
@@
For unconstrained decoding methods, we run Nucleus Sampling and Beam Sampling. To have a fair comparison against results in SVEN, we follow the procedure outlined in the paper (He and Vechev, 2023). For each model, we generate 25 code completions given each prompt. We run the experiment 10 times using different random seeds. We calculate the performance metrics for each experiment. Then, we present the average results across the experiments, as well as the 95% confidence intervals.
Table 2. Performance (%) of different decoding methods over CodeGen and SVEN. We report the metrics over CodeGuard+, including the mean values of 10 random seeds and the 95% confidence intervals in the parentheses. The best number in each column is highlighted in bold. Our new metric secure-pass@$1$ is more realistic than SVEN-SR, since we evaluate both security and correctness of generated code, but SVEN-SR ignores correctness. Constrained beam sampling improves secure-pass@$1$ for both CodeGen and SVEN.
Model Decoding Method pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR Constraint Rate
CodeGen Nucleus 66.66 (±$\pm 1.28$) 59.50 (±$\pm 2.00$) 43.17 (±$\pm 1.39$) 60.65 (±$\pm 0.78$) N/A
Beam 86.17 (±$\pm 2.09$) 66.89 (±$\pm 3.90$) 63.74 (±$\pm 2.48$) 71.19 (±$\pm 2.84$) N/A
Constrained Beam 76.00 (±$\pm 0.24$) 79.13 (±$\pm 2.41$) 76.00 (±$\pm 0.24$) 91.30 (±$\pm 0.00$) 74.81 (±$\pm 0.35$)
SVEN Nucleus 55.60 (±$\pm 1.11$) 80.01 (±$\pm 2.76$) 47.48 (±$\pm 1.25$) 84.49 (±$\pm 0.76$) N/A
Beam 71.63 (±$\pm 2.36$) 69.44 (±$\pm 3.55$) 62.19 (±$\pm 3.12$) 84.01 (±$\pm 0.83$) N/A
Constrained Beam 82.17 (±$\pm 1.01$) 83.48 (±$\pm 2.41$) 82.17 (±$\pm 1.01$) 87.83 (±$\pm 2.41$) 70.03 (±$\pm 4.25$)
@@
For constrained decoding methods, we run our Constrained Beam Sampling and our adapted MuCoLa. We evaluate them in a setting where we want all outputs to satisfy the constraints. For each prompt, we generate 10 completions that satisfy constraints. Since sometimes the method may not generate an output that satisfies the constraints, we continue the generation until we get 10 constrained outputs, or until we reach a maximum of 100 outputs, whichever happens first. Then, we repeat this experiment five times with five different seeds. We calculate the performance metrics for each experiment. If no generation meets the constraints, we assign a value of 0 to all metrics. Then, we present the average results across experiments, as well as the 95% confidence intervals.
We evaluate all decoding methods over CodeGen and SVEN. The details of the hyperparameters can be found in Appendix C. Since MuCoLa can only work with models with the same input and output embedding layers, we only evaluate MuCoLa over StarCoder2-3B. CodeGen-2.7B does not share weights between its input and output embedding layers. Previously, MuCoLa is only tested on GPT-2 family models. We discuss engineering lessons to make MuCoLa work on StarCoder2 in Appendix E. We run all experiments on a cluster with NVIDIA A100 GPUs (80 GB).
5.2. Performance of Unconstrained Decoding
@@
We explore whether using different decoding methods changes how a Code LLM generates secure and correct code. We compare the performance of Nucleus Sampling and Beam Sampling over CodeGen and SVEN, with results in Table 2. For both models, Beam Sampling outperforms Nucleus Sampling. For CodeGen, Beam Sampling has 19.51% higher pass@$1$ and 20.57% higher secure-pass@$1$ than Nucleus Sampling. For SVEN, Beam Sampling has 16.03% higher pass@$1$ and 14.71% higher secure-pass@$1$ than Nucleus Sampling, even though secure@$1_{\text{pass}}$ decreases by 10.57%. The results show that Beam Sampling makes the models more likely to generate correct and secure code.
Key Result: Different decoding methods make a big difference in the quality of generated code, in terms of security and functional correctness. For CodeGen, Beam Sampling has 20.57% higher secure-pass@$1$ than Nucleus Sampling.
@@
Across all settings in Table 2, SVEN-SR is much higher than secure-pass@$1$. This is mainly due to the fact that SVEN-SR only evaluates whether the generated code is secure, ignoring whether they are also correct. The big drop from SVEN-SR to secure-pass@$1$ can be explained by the values of pass@$1$. For example, when running Nucleus Sampling over SVEN, secure-pass@$1$ is only 55.60%. This means that almost half of the generated code is wrong. Since SVEN-SR is 84.49% in this setting, whereas secure-pass@$1$ is only 47.48%, this may be interpreted as, about half of the secure code is incorrect. We see similar trends in other settings that lower pass@$1$ correlates with lower secure-pass@$1$, but higher pass@$1$ correlates with higher secure-pass@$1$. For example, Nucleus Sampling and Beam Sampling over SVEN have almost the same SVEN-SR (both 84%), but Beam Sampling has a much higher pass@$1$ than Nucleus Sampling, which makes the secure-pass@$1$ for Beam Sampling higher too.
Key Result: SVEN-SR severely overestimates the security level of Code LLMs, overlooking whether the generated secure code is correct. Our new metric secure-pass@$1$ is a more realistic measure of security and correctness of Code LLMs.
@@
First, we compare CodeGen with SVEN using Nucleus sampling, the same setting in the SVEN paper (He and Vechev, 2023). The secure-pass@$1$ of SVEN is 47.48%, only 4.31% higher than CodeGen. Second, when we use Beam Sampling, SVEN does not have any advantage in secure-pass@$1$. CodeGen has 63.74% secure-pass@$1$, even 1.55% higher than SVEN.
We also notice the tension between security and functional correctness in SVEN. SVEN increases secure@$1_{\text{pass}}$ by 20.51% compared to CodeGen when using Nucleus Sampling, meaning it increases the likelihood of generating secure code when the code is correct. However, it also decreases pass@$1$ by 11.06% compared to CodeGen. Consequently, the advantage of SVEN to generate code that is both secure and correct was not as strong as previously thought, and this advantage disappears under Beam Sampling.
Figure 7. We list 8 prompts where the conclusion of comparing SVEN to CodeGen has reversed. In these test scenarios, SVEN has higher (or equivalent) SVEN-SR than CodeGen, but lower secure-pass@$1$ than CodeGen, when using Nuclues Sampling.
Key Result: SVEN achieves a 4.31% improvement of secure-pass@$1$ over CodeGen when using Nucleus Sampling, and SVEN is no better than CodeGen with Beam Sampling. SVEN improves security by sacrificing functional correctness.
@@
When using Nucleus Sampling, SVEN has worse secure-pass@$1$ than CodeGen in 8 prompts, even though SVEN has higher SVEN-SR than CodeGen for these prompts. We plot them in Figure 7. From CodeGen to SVEN, the decrease in secure-pass@$1$ ranges from 1.2% (for “CWE-125 2-c”) to 72.8% (for “CWE-089 0-py”). For “CWE-079 0-py”, SVEN achieves 100% SVEN-SR, compared to CodeGen’s 35.2%; but the secure-pass@$1$ score of SVEN is only 4%, compared to CodeGen’s 26.4%. We present all breakdown results for CodeGen and SVEN in Table 6 in Appendix D. One example of safe but incorrect generation of SVEN is shown in Figure 8. We find that SVEN is more likely to generate incomplete SQL queries compared to CodeGen in this case.
@@
# create a cursor to execute queries on the database
@@
Figure 8. An example secure generation by SVEN over “CWE-089 0-py”. The generated content is highlighted. There is an incomplete SQL query "SELECT?". SVEN is more likely to generate incomplete SQL queries compared to CodeGen.
Table 3. Performance (%) of different decoding schemes over StarCoder2. Constrained Beam Sampling has a similar secure-pass@$1$ as Beam Sampling, due to the low Constraint Rate. We find that MuCoLa struggles with generating correct code (low pass@$1$), and it does not perform well on code generation.
Model Decoding Method pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR Constraint Rate
StarCoder2 Nucleus 81.50 (±$\pm 0.80$) 66.55 (±$\pm 1.48$) 52.49 (±$\pm 0.84$) 66.46 (±$\pm 1.13$) N/A
Beam 85.83 (±$\pm 1.23$) 66.80 (±$\pm 2.28$) 65.69 (±$\pm 1.76$) 75.57 (±$\pm 2.37$) N/A
Constrained Beam 66.17 (±$\pm 3.47$) 66.96 (±$\pm 2.96$) 65.30 (±$\pm 1.45$) 69.57 (±$\pm 0.00$) 57.66 (±$\pm 3.76$)
MuCoLa 51.66 (±$\pm 3.89$) 82.37 (±$\pm 6.67$) 46.06 (±$\pm 2.96$) 88.86 (±$\pm 4.85$) 50.29 (±$\pm 2.17$)
Table 4. Performance (%) of different decoding schemes over StarCoder2 on a subset of CodeGuard+ where the model has high Constraint Rate. Constrained Beam Sampling achieves the best performance for secure@$1_{\text{pass}}$ and secure-pass@$1$. This suggests that improving the Constraint Rate may improve the performance of Constrained Beam Sampling.
Model # Prompt Decoding Method pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR Constraint Rate
StarCoder2 17$\dagger$ Nucleus 82.73 (±$\pm 0.62$) 75.38 (±$\pm 1.71$) 62.14 (±$\pm 0.63$) 74.67 (±$\pm 1.40$) N/A
Beam 92.56 (±$\pm 1.74$) 83.91 (±$\pm 3.30$) 82.94 (±$\pm 2.43$) 84.60 (±$\pm 3.21$) N/A
Constrained Beam 88.35 (±$\pm 1.96$) 90.59 (±$\pm 4.00$) 88.35 (±$\pm 1.96$) 94.12 (±$\pm 0.00$) 77.99 (±$\pm 5.07$)
  • $\dagger$
    6 test cases, “CWE-022 1-py”, “CWE-022 2-py”, “CWE-078 2-py”, “CWE-078 2-py”, “CWE-476 0-c”, “CWE-476 2-c” and “CWE-787 0-c”, are removed.
5.3. Performance of Constrained Decoding
Constrained Decoding on CodeGen and SVEN

### v2

@@ 4.4 Non-autoregressive Decoding
Non-autoregressive decoding generates all tokens in the output sequence together. The decoding procedure first initializes output tokens, and then uses gradients of some function to update the tokens. An example function could be a language model loss function, an energy function, or some task-specific function. Non-autoregressive decoding methods have shown promising results for machine translation [43], reasoning and counterfactual story generation [44], and generative commonsense reasoning [39, 14].
@@
Recent papers argue that non-autoregressive decoding is better than autoregressive decoding for the problem of controlled text generation under constraints [39, 44, 14]. The same arguments hold for code generation under constraints. During autoregressive decoding, we cannot evaluate the properties of the entire program during the generation because only a partial program is available at every step. For example, if the partially generated code has not sanitized untrusted user input yet, it does not mean that the entire generated code would not sanitize untrusted user input, so we cannot know whether the partial program is safe or not safe. On the contrary, non-autoregressive decoding generates the entire program altogether, which enables us to evaluate constraints as well as enforce constraints over the whole program.
@@
To the best of our knowledge, non-autoregressive decoding has not been evaluated on code generation before, but only text generation. In particular, the state-of-the-art scheme MuCoLa [14] has achieved strong results of constrained text generation for common sense reasoning, beating previous methods. Therefore, we study MuCoLa and adapt it for code generation.
@@
$\mathbf{y}\sim P(\mathbf{y}|\mathbf{x}),$ (10)
@@
MuCoLa formulates decoding as sampling from an energy-based model (EBM) using Langevin Dynamics, following the approach in COLD decoding [44]. In other words, MuCoLa performs sampling by iteratively updating the embeddings of the output sequence using gradients of the energy function. They define the energy function as the following:
$\mathcal{E}(\tilde{\mathbf{e}})=-\log P(\tilde{\mathbf{e}}|\mathbf{x})-\sum_{i=1}^{C}\lambda_{i}\left(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\right).$ (11)
Here, $\lambda_{i}$ is used to balance between the output fluency and satisfying constraints. MuCoLa uses gradients to perform sampling, with details in Section .2. The gradient update procedure will converge to sampling from the energy-based distribution [45].
@@
We adapt MuCoLa for constrained code generation using our security constraints. We can separate our constraints into positive constraints and negative constraints. Positive constraints are key phrases that we would like to appear in generated outputs, and negative constraints are key phrases we want to block, where each phrase consists of multiple tokens. We have in total $C^{+}$ positive constraints, and $C^{-}$ negative constraints.
MuCoLa provides a differentiable positive key phrase function $f$ (details in Section .2). We use that as a building block to define our own energy function:
$\mathcal{E}^{\prime}(\tilde{\mathbf{e}})=$ $-\log P(\tilde{\mathbf{e}}|\mathbf{x})-\sum_{i=1}^{C^{+}}\lambda_{i}\left(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\right)$ (12)
$-\sum_{j=1}^{C^{-}}\lambda_{j}\left(f_{j}(\tilde{\mathbf{e}})-\epsilon_{j}\right).$
@@ 5 Evaluation
  • RQ2. If we use our new metrics to compare a baseline Code LLM against the state-of-the-art prefix tuning defense SVEN [7], how does that change the conclusions about the defense? (Section 5.2)
  • RQ3. Can constrained decoding improve the security and correctness of code generated by Code LLMs? (Section 5.3)
@@ 5.1 Experiment Setup
We evaluate eight state-of-the-art (SOTA) Code LLMs in total, ordered by the model size: CodeGen-2.7B [18], SVEN [7], StarCoder2-3B [19], CodeGemma-7B [20], Llama3-8B [21], DeepseekCoder-33B [22], CodeLlama-34B [23] and GPT-4 [17]. The first seven models are SOTA open-source, decoder-only pre-trained models, and the last model is the proprietary GPT-4 model from OpenAI. Among them, SVEN is the secure CodeGen-2.7B with the SOTA prefix tuning defense. We use the trained prefix on CodeGen-2.7B to generate secure code released by the authors of SVEN [7].
@@
We use the CodeGuard+ introduced in Section 3.1 as the test suite for all evaluations. This suite contains 91 prompts covering 34 CWEs and 2 programming languages C/C++ and Python. We use our unit tests to evaluate correctness. Related works [7, 6, 8, 13] use either CodeQL [46] or Sonar [47] to automatically evaluate the security of generated code. In our experiments, we use an ensemble of both CodeQL and Sonar: if any of the two static analyzers say the generated code is vulnerable, we detect that as vulnerable; only if both static analyzers consider the generated code as secure, we evaluate that as secure. We present our evaluation results using four metrics: SVEN-SR, pass@$1$, secure@$1_{\text{pass}}$, and secure-pass@$1$, which are defined in Section 3.2.
@@
For unconstrained decoding over open-source models, we run Nucleus Sampling and Beam Sampling. For each open-source model, we generate 10 code completions given each prompt. We run the experiment 10 times using different random seeds. We calculate the performance metrics for each experiment. Then, we present the average results across the experiments, as well as the 95% confidence intervals.
For unconstrained decoding over GPT-4, we run Nucleus Sampling. We generate 25 code completions given each prompt. We describe the prompt templates for querying GPT-4 and the post-processing procedure for GPT-4’s generations in  Section .3.
@@
For constrained decoding methods, we run our Constrained Beam Sampling and our adapted MuCoLa in a setting where we want all outputs to satisfy the constraints. For each prompt, we generate 10 completions that satisfy constraints. Since sometimes the method may not generate an output that satisfies the constraints, we continue the generation until we get 10 constrained outputs, or until we reach a maximum number of outputs, whichever happens first. For Constrained Beam Sampling, we use a maximum of 100 outputs per prompt and repeat this experiment 10 times with different seeds; for MuCoLa, we use a maximum of 30 outputs per prompt and repeat this experiment 5 times with different seeds, since MuCoLa runs relatively slower. We calculate the performance metrics for each experiment. If no generation meets the constraints, we assign 0 to all metrics. Then, we present the average results across experiments, as well as the 95% confidence intervals. We evaluate Constrained Beam Sampling over all open-source models. Since MuCoLa can only work with models with the same input and output embedding layers, we only evaluate MuCoLa over StarCoder2-3B. We discuss engineering lessons to make MuCoLa work on StarCoder2 in Section .5.
The details of the hyperparameters can be found in Section .4. We run experiments on a cluster with NVIDIA A100 GPUs (80 GB) as well as on a server with four NVIDIA H100 GPUs (80 GB).
TABLE III: Performance of different decoding schemes over Code LLMs, evaluated over CodeGuard+. We report mean values (%) across different random seeds and the 95% confidence intervals in the parentheses. The best number in each column is highlighted in bold. Our new metric secure-pass@$1$ is more realistic than SVEN-SR, since we evaluate both security and correctness of generated code, but SVEN-SR ignores correctness. Constrained Beam Sampling for all open-source models except SVEN outperforms GPT-4 with unconstrained decoding in secure-pass@$1$.
Model Decoding Method pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR
CodeGen-2.7B Nucleus 49.89 (±$\pm 0.63$) 40.86 (±$\pm 2.15$) 26.07 (±$\pm 0.81$) 53.65 (±$\pm 1.64$)
Beam 61.35 (±$\pm 0.92$) 37.26 (±$\pm 1.36$) 33.77 (±$\pm 1.23$) 54.08 (±$\pm 1.87$)
Constrained Beam 61.05 (±$\pm 1.01$) 56.04 (±$\pm 1.17$) 51.25 (±$\pm 1.17$) 79.12 (±$\pm 1.69$)
SVEN Nucleus 42.95 (±$\pm 0.95$) 51.80 (±$\pm 1.81$) 29.14 (±$\pm 0.78$) 71.91 (±$\pm 0.95$)
Beam 55.24 (±$\pm 1.22$) 40.84 (±$\pm 1.09$) 36.30 (±$\pm 1.29$) 67.42 (±$\pm 1.50$)
Constrained Beam 54.20 (±$\pm 0.55$) 49.76 (±$\pm 0.84$) 46.26 (±$\pm 0.69$) 79.38 (±$\pm 2.26$)
StarCoder2-3B Nucleus 70.80 (±$\pm 0.49$) 52.13 (±$\pm 1.15$) 38.88 (±$\pm 0.55$) 55.43 (±$\pm 0.74$)
Beam 77.62 (±$\pm 1.03$) 47.89 (±$\pm 1.49$) 46.12 (±$\pm 1.51$) 55.80 (±$\pm 1.12$)
Constrained Beam 70.79 (±$\pm 1.12$) 59.76 (±$\pm 1.17$) 59.56 (±$\pm 1.31$) 70.79 (±$\pm 1.78$)
MuCoLa 46.07 (±$\pm 2.43$) 66.66 (±$\pm 2.27$) 39.60 (±$\pm 1.62$) 80.17 (±$\pm 1.47$)
CodeGemma-7B Nucleus 73.93 (±$\pm 0.67$) 54.34 (±$\pm 1.21$) 43.63 (±$\pm 0.83$) 59.85 (±$\pm 0.92$)
Beam 78.41 (±$\pm 1.13$) 51.88 (±$\pm 0.55$) 50.46 (±$\pm 0.62$) 61.86 (±$\pm 1.01$)
Constrained Beam 76.22 (±$\pm 1.41$) 61.37 (±$\pm 1.18$) 59.34 (±$\pm 1.29$) 74.22 (±$\pm 0.91$)
Llama3-8B Nucleus 74.37 (±$\pm 0.58$) 57.88 (±$\pm 0.88$) 46.54 (±$\pm 0.64$) 61.61 (±$\pm 0.52$)
Beam 82.20 (±$\pm 1.09$) 50.41 (±$\pm 1.15$) 49.93 (±$\pm 1.07$) 58.04 (±$\pm 1.13$)
Constrained Beam 72.21 (±$\pm 0.71$) 61.41 (±$\pm 0.77$) 58.48 (±$\pm 0.60$) 70.76 (±$\pm 1.08$)
DeepseekCoder-33B Nucleus 78.77 (±$\pm 0.84$) 56.09 (±$\pm 0.93$) 46.54 (±$\pm 1.11$) 60.12 (±$\pm 0.71$)
Beam 80.65 (±$\pm 0.71$) 47.37 (±$\pm 1.71$) 46.58 (±$\pm 1.51$) 57.25 (±$\pm 1.06$)
Constrained Beam 71.87 (±$\pm 0.55$) 60.15 (±$\pm 1.49$) 57.97 (±$\pm 1.56$) 77.57 (±$\pm 1.83$)
CodeLlama-34B Nucleus 75.47 (±$\pm 0.64$) 53.51 (±$\pm 0.57$) 44.53 (±$\pm 0.32$) 55.36 (±$\pm 0.58$)
Beam 81.65 (±$\pm 0.94$) 49.73 (±$\pm 0.60$) 48.99 (±$\pm 0.70$) 54.24 (±$\pm 0.50$)
Constrained Beam 69.84 (±$\pm 0.77$) 63.31 (±$\pm 1.22$) 60.85 (±$\pm 0.88$) 72.37 (±$\pm 0.93$)
GPT-4 Nucleus 70.13 57.97 47.45 63.67
5.2 Performance of Unconstrained Decoding
@@
We explore whether using different decoding methods changes how a Code LLM generates secure and correct code. We compare the performance of Nucleus Sampling and Beam Sampling over CodeGen, SVEN, StarCoder2, CodeGemma, Llama3, DeepseekCoder, and CodeLlama, with results in Table III. The results show that Beam Sampling makes the models more likely to generate correct and secure code. For CodeGen, Beam Sampling has 11.46% higher pass@$1$ and 7.7% higher secure-pass@$1$ than Nucleus Sampling. For SVEN, Beam Sampling has 12.29% higher pass@$1$ and 7.16% higher secure-pass@$1$ than Nucleus Sampling, even though secure@$1_{\text{pass}}$ decreases by 10.96%. We observe similar trends for StarCoder2, CodeGemma, Llama3, and CodeLlama. Only for DeepseekCoder, Beam Sampling has a similar secure-pass@$1$ as Nucleus Sampling.
Key Result: Different decoding methods make a big difference in the quality of generated code, in terms of security and functional correctness. For six open-source models, Beam Sampling has higher pass@$1$ and higher secure-pass@$1$ than Nucleus Sampling.
@@
Across all settings in Table III, SVEN-SR is much higher than secure-pass@$1$. This is mainly due to the fact that SVEN-SR only evaluates whether the generated code is secure, ignoring whether they are also correct. The big drop from SVEN-SR to secure-pass@$1$ can be explained by the values of pass@$1$. For example, when running Nucleus Sampling over SVEN, SVEN-SR is 71.91%, whereas secure-pass@$1$ is only 29.14%. This may be interpreted as, the majority of generated secure code is incorrect. We see similar trends in other settings that lower pass@$1$ correlates with lower secure-pass@$1$, but higher pass@$1$ correlates with higher secure-pass@$1$. For example, Nucleus Sampling and Beam Sampling over SVEN have similar SVEN-SR (71.91% vs 67.42%), but Beam Sampling has a much higher pass@$1$ than Nucleus Sampling, which makes the secure-pass@$1$ for Beam Sampling higher too.
Key Result: SVEN-SR severely overestimates the security level of Code LLMs, overlooking whether the generated secure code is correct. Our new metric secure-pass@$1$ is a more realistic measure of the security of Code LLMs.
@@
First, we compare CodeGen with SVEN using Nucleus sampling, the same setting in the SVEN paper [7]. The secure-pass@$1$ of SVEN is 29.14%, only 3.07% higher than CodeGen. Second, when we use Beam Sampling, SVEN has 36.3% secure-pass@$1$, only 2.53% higher than SVEN.
We also notice the tension between security and functional correctness in SVEN. SVEN increases secure@$1_{\text{pass}}$ by 10.94% compared to CodeGen when using Nucleus Sampling, meaning it increases the likelihood of generating secure code when the code is correct. However, it also decreases pass@$1$ by 6.94% compared to CodeGen. Consequently, the advantage of SVEN to generate code that is both secure and correct was not as strong as previously thought.
Figure 7: We list 11 prompts where the conclusion of comparing SVEN to CodeGen has reversed. In these test scenarios, SVEN has higher (or equivalent) SVEN-SR than CodeGen, but lower secure-pass@$1$ than CodeGen, using Nuclues Sampling.
Key Result: SVEN achieves only 3.07% improvement of secure-pass@$1$ over CodeGen when using Nucleus Sampling, and only 2.53% improvement with Beam Sampling. SVEN improves security by sacrificing functional correctness.
@@
SVEN has used SVEN-SR to show superior performance over prompts in the 9 CWEs they cover in the training set [7]. Thus, we use our new metric secure-pass@$1$ to study the performance of the models using 33 prompts belonging to these 9 CWEs in CodeGuard+. When using Nucleus Sampling, SVEN has worse secure-pass@$1$ than CodeGen in 11 prompts, even though SVEN has higher (or equivalent) SVEN-SR than CodeGen for these prompts, as shown in Figure 7. From CodeGen to SVEN, the decrease in secure-pass@$1$ ranges from 1% (for “CWE-125 0-c”) to 74% (for “CWE-089 0-py”). For “CWE-079 0-py”, SVEN achieves 100% SVEN-SR, compared to CodeGen’s 22.8%. However, the secure-pass@$1$ score of SVEN is only 4%, compared to CodeGen’s 18%. One example of safe but incorrect generation of SVEN is shown in Figure 8. We find that SVEN is more likely to generate incomplete SQL queries compared to CodeGen in this case.
@@
# create a cursor to execute queries on the
# database
@@
Figure 8: An example of secure but incorrect generation by SVEN over “CWE-089 0-py”. The generated content is highlighted. There is an incomplete SQL query "SELECT?".
5.3 Performance of Constrained Decoding
Constrained Decoding vs Prefix Tuning